# Alignment of Diffusion Models: Fundamentals, Challenges, and Future

### v3

@@ Abstract.
Diffusion models have emerged as the leading paradigm in generative modeling, excelling in various applications. Despite their success, these models often misalign with human intentions and generate results with undesired properties or even harmful content. Inspired by the success and popularity of alignment in tuning large language models, recent studies have investigated aligning diffusion models with human expectations and preferences. This work mainly reviews alignment of text-to-image diffusion models, covering advancements in fundamentals of alignment, alignment techniques of diffusion models, preference benchmarks, and evaluation for diffusion models. Moreover, we discuss key perspectives on current challenges and promising future directions on solving the remaining challenges in alignment of diffusion models. To the best of our knowledge, our work is the first comprehensive review paper for researchers and engineers to comprehend, practice, and research alignment of diffusion models.  [^1] https://github.com/xie-lab-ml/awesome-alignment-of-diffusion-models
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language generation
@@ 1. Introduction
[…]
The left pie chart illustrates the distribution of total papers, showing that LLMs account for 84.5% of the research (82,000 papers), compared to 15.5% (15,000 papers) for diffusion models. The right pie chart highlights alignment research, revealing a significant skew towards LLMs, with 98.0% of alignment-related studies (18,300 papers) focused on LLMs, while only 2.0% (374 papers) address alignment in diffusion models. This disparity underscores the relatively nascent focus on alignment within diffusion models compared to the research on LLMs.
Figure 2. Statistical overview of research trends. (a) The number of papers on diffusion models at top computer vision conferences (CVPR, ECCV, ICCV) and top machine learning conferences (NeurIPS, ICML, ICLR) since 2017, indicating a growing interest. Note that NeurIPS 2025 has not been released as of the date of submission, and ECCV and ICCV are held biennially. (b) Comparison of the ratio of papers on LLMs vs. diffusion models (left pie) and the research focus on alignment within these areas (right pie), highlighting the nascent stage of diffusion model alignment.
Inspired by the success of aligning LLMs (Wang et al., 2023d), there is growing interest in better aligning diffusion models with human intentions to enhance their capabilities. Fig. 2(b) visualizes the paper counts on LLMs and diffusion models, as well as their alignment studies  [^2] Data obtained from Google Scholar as of July 25, 2025. . The left pie chart shows LLMs account for 84.5% of the studies, while diffusion models account for 15.5%. The right chart highlights that, at this point, 98.0% of alignment studies focus on LLMs, while only 2.0% address diffusion models. This disparity, clearly illustrated in Fig. 2(b), not only underscores the relatively nascent stage of alignment research for diffusion models compared to LLMs but also signals a significant opportunity and pressing need for focused investigation in this domain. This survey aims to facilitate such research by consolidating current knowledge and outlining future directions. The fundamental differences between LLMs and diffusion models, where LLMs predict sequences of tokens and diffusion models progressively reverse a noise-adding diffusion process, and where the continuous, high-dimensional nature of image generation versus discrete token prediction in LLMs presents unique challenges for defining preferences and applying feedback, along with their uniquely advantageous application domains, such as LLMs for language generation and diffusion models for image generation, make the study of diffusion model alignment an independent and valuable area of interest.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=10em,font=,, where level=3text width=11em,font=,, where level=4text width=7em,font=,, where level=5text width=7em,font=,, [ Alignment of Diffusion Models, ver [ Fundamentals of
Human Alignment [ Preference Data
and Modeling [ Bradley-Terry Model  (Bradley and Terry, 1952),
Plackett-Luce Model  (Plackett, 1975; Luce, 1959), leaf, text width=18em ] ] [ Alignment
Algorithms [ Reinforcement Learning
from Human Feedback [ InstructGPT  (Ouyang et al., 2022), REINFORCE  (Ahmadian et al., 2024),
RAFT  (Dong et al., 2023), RRHF  (Yuan et al., 2023), RSO  (Liu et al., 2024c),
CAI  (Bai et al., 2022), RLAIF  (Lee et al., 2024a), leaf, text width=15em ] ] [ Direct Preference
Optimization [ DPO  (Rafailov et al., 2023), IPO  (Gheshlaghi Azar et al., 2024), ORPO  (Hong et al., 2024a),
KTO  (Ethayarajh et al., 2024), PRO  (Song et al., 2024), LiPO  (Liu et al., 2024a), leaf, text width=15em ] ] ] ] [ Human Alignment of
Diffusion Models [ Training-based
Alignment [ RLHF [ Lee et al. (2023a), ReFL  (Xu et al., 2023), DPOK  (Fan et al., 2023),
DDPO  (Black et al., 2023), PRDP  (Deng et al., 2024), DRaFT  (Clark et al., 2024),
AlignProp  (Prabhudesai et al., 2023), DRTune  (Wu et al., 2024b) , leaf, text width=18em ] ] [ DPO [ Diffusion-DPO  (Wallace et al., 2024), D3PO  (Yang et al., 2024d),
Dense Reward (Yang et al., 2024a), SPO (Liang et al., 2024b),
Diffusion-KTO (Li et al., 2024a) , leaf, text width=18em ] ] ] [ Test-time Alignment [ Implicit Guidance [ Prompt Optimization, leaf, text width=12em ] [ Attention Control, leaf, text width=12em ] [ Initial Noise Optimization, leaf, text width=12em ] ] [ Explicit Guidance [ Reward-based
Input Optimization, leaf, text width=12em ] [ Reward-Guided
Decoding and Sampling, leaf, text width=12em ] ] ] [ Alignment Beyond
T2I Diffusion Models [ Video Generation (Prabhudesai et al., 2024; Yuan et al., 2024b; Wang et al., 2024b; Xu et al., 2024b; Liu et al., 2025; Zhang et al., 2024c; Jiang et al., 2025; Huang et al., 2025a; Lee et al., 2024b; Oshima et al., 2025),
Image Editing  (Zhang et al., 2024d), 3D Generation  (Ye et al., 2024; Zhou et al., 2025b; Ignatyev et al., 2025),
Molecule Generation  (Gu et al., 2024), Decision Making  (Dong et al., 2024b),
Audio Generation (Majumder et al., 2024), Motion Generation (Tan et al., 2024), leaf, text width=30em ] ] ] [ Benchmarks and
Evaluation [ Benchmarks [ Scalar Human
Preference Dataset [ HPD v1 (Wu et al., 2023b), HPD v2 (Wu et al., 2023a),
Pick-a-Pic v1 (Kirstain et al., 2023), ImageRewardDB  (Xu et al., 2023),
Picsart Image-Social  (Isajanyan et al., 2024), leaf, text width=18em ] ] [ Multi-dimensional Human
Feedback Dataset [ MHP (Zhang et al., 2024b), RichHF-18K (Liang et al., 2024a), leaf, text width=18em ] ] [ AI Feedback Dataset [ VisionPrefer (Wu et al., 2024c), DreamBench++ (Peng et al., 2025), leaf, text width=18em ] ] ] [ Evaluation [ Reward Models [ Human Feedback Prediction,
Correlation with Human Raters, leaf, text width=18em ] ] [ T2I Diffusion Models [ Prompts [ MS-COCO (Lin et al., 2014), DiffusionDB (Wang et al., 2023b),
MT-Bench (Petsiuk et al., 2022), leaf, text width=16em ] ] [ Image Quality [ IS (Salimans et al., 2016), FID (Heusel et al., 2017), leaf, text width=16em ] ] [ Human
Preference
Evaluation [ CLIP (Radford et al., 2021), Aesthetic (Schuhmann et al., 2022),
HPS v1 (Wu et al., 2023b), HPS v2 (Wu et al., 2023a),
PickScore (Kirstain et al., 2023), ImageReward  (Xu et al., 2023),
MPS  (Zhang et al., 2024b), VP-Score (Wu et al., 2024c),
Social Reward  (Isajanyan et al., 2024),leaf, text width=16em ] ] [ Fine-grained
Evaluation [ DALL-Eval (Cho et al., 2023a), GenEval (Ghosh et al., 2023),
VPEval (Cho et al., 2023b), HEIM (Lee et al., 2023b),
LLMScore (Lu et al., 2023), Style (Somepalli et al., 2024), leaf, text width=16em ] ] ] ] ] ]
@@ 3.1. Preference Data and Modeling
Preference data are typically composed of prompts, responses, and feedback. Prompts, whether human-provided or AI-generated, are diverse inputs used to elicit model responses. Responses can be generated on-policy (from the current model) or off-policy (from external or older models), presenting a trade-off between relevance and collection efficiency (Sutton and Barto, 2020).
@@ 4. Human Alignment Techniques of Diffusion Models
RLHF and DPO are two very classic training-based techniques for aligning AI models with human preferences. However, when applied to diffusion models, these methods encounter significant challenges due to the step-by-step training and sampling nature of diffusion models. Specifically, aligning diffusion models with preference optimization requires sampling across all possible diffusion trajectories leading to $x_{0}$, which is intractable in practice. While the LLM response is treated as a single output, diffusion models’ multiple latent image representations of each step need to be calculated and stored, leading to unreasonable high memory consumption and low computation efficiency. This makes these methods impractical for large-scale diffusion models. To address the high computational overhead associated with adapting alignment techniques to diffusion models, researchers often formulate the denoising process as a multi-step Markov decision process (MDP). The proposed diffusion alignment methods need to directly optimize the expected reward of an image output or update the policy based on human preferences to approximately perform policy gradient guided by a reward model. This formulation enables parameter updates at each step of the denoising process based on human preferences, thereby circumventing the significant computational costs.
@@ 4.3.3. Summary and Outlook
Test-time alignment techniques provide efficient and flexible pathways to improve generation without retraining. Their evolution clearly shows a trend from implicit guidance to explicit reward-guided strategies.
Implicit methods, such as prompt optimization and attention control, are simple to implement and add little inference overhead, making them effective for lightweight alignment. However, their control is less precise and often relies on heuristics. In contrast, explicit reward-guided methods offer more powerful and direct control by incorporating external preference models, enabling them to handle more complex alignment tasks. Their main challenges are the increased computational cost at inference time and the potential for “reward hacking”, where the model optimizes the proxy score rather than the true image quality.
@@ 4.5. Open Challenges and Research Frontiers in Diffusion Alignment
In this subsection, we discuss several open challenges and research frontiers in diffusion alignment, focusing on the nuanced comparison between dominant paradigms, insights from LLM alignment, and challenges unique to the diffusion models.
@@ 5.1.3. AI Feedback Datasets
HPD v1 $\rightarrow$ HPS v1
 (Wu et al., 2023b)
Stable Foundation
Discord channel
25,205 Stable Diffusion 98,807
2659
experienced users
25,205
HPD v2 $\rightarrow$ HPS v2
 (Wu et al., 2023a)
COCO Captions +
DiffusionDB
107,915
9 models +
COCO images
433,760
57
employed annotators
798,090
Pick-a-Pic v1 $\rightarrow$ PickScore
 (Kirstain et al., 2023)
Real users 37,523
Stable Diffusion
variants
1,169,494
6,394
web app users
584,747
ImageRewardDB $\rightarrow$ ImageReward
 (Xu et al., 2023)
DiffusionDB 8,878 6 models 273,784
Annotation
company
136,892
MHP $\rightarrow$ MPS
 (Zhang et al., 2024b)
PromptHero +
DiffusionDB +
KOLORS + GPT4
66,389 9 models 607,541
210 crowd-sourced
annotators
918,315
RichHF-18K $\rightarrow$ RAHF
 (Liang et al., 2024a)
Pick-a-Pic v1 17,760 Pick-a-Pic v1 35,520
27
trained annotators
17,760
Picsart Image-Social $\rightarrow$ Social Reward
 (Isajanyan et al., 2024)
Social platform
user prompts
104 K
Several
in-house models
1.7 M
1.5 M
users
3M
VisionPrefer $\rightarrow$ VP-Score
 (Wu et al., 2024c)
DiffusionDB
179 K
Stable Diffusion
variants
0.76 M
GPT-4 Vision
1.2 M
@@ 5.2.1. Evaluation for Reward Models
CLIP score (Radford et al., 2021)
54.3 71.2 63.7 60.8* 51.9
Aesthetic score (Schuhmann et al., 2022)
57.4 72.6 62.9 56.8* 55.3
HPS v1 (Wu et al., 2023b)
61.2 73.1 65.5 66.7* -
HPS v2 (Wu et al., 2023a)
65.7 83.3 65.5 67.4* 59.4
PickScore (Kirstain et al., 2023)
62.9 79.8 69.5 70.5* 62.6
ImageReward (Xu et al., 2023)
65.1 70.6 67.5 61.1* 60.5
MPS (Zhang et al., 2024b)
67.5 83.5 74.2 - -
VP-Score (Wu et al., 2024c)
66.3* 79.4* - 67.1* -
Social Reward (Isajanyan et al., 2024)
- - - - 69.7

### v4

@@ Abstract.
Diffusion models have emerged as the leading paradigm in generative modeling, excelling in various applications. Despite their success, these models often misalign with human intentions and generate results with undesired properties or even harmful content. Inspired by the success and popularity of alignment in tuning large language models, recent studies have investigated aligning diffusion models with human expectations and preferences. This work mainly reviews alignment of diffusion models, covering advancements in fundamentals of alignment, alignment techniques of diffusion models, preference benchmarks, and evaluation for diffusion models. Moreover, we discuss key perspectives on current challenges and promising future directions on solving the remaining challenges in alignment of diffusion models. To the best of our knowledge, our work is the first comprehensive review paper for researchers and engineers to comprehend, practice, and research alignment of diffusion models.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language generation
@@ 1. Introduction
The left pie chart illustrates the distribution of total papers, showing that LLMs account for 89.4% of the research (209,000 papers), compared to 10.6% (24,700 papers) for diffusion models. The right pie chart highlights alignment research, revealing a significant skew towards LLMs, with 93.5% of alignment-related studies (8,300 papers) focused on LLMs, while only 6.5% (573 papers) address alignment in diffusion models. This disparity underscores the relatively nascent focus on alignment within diffusion models compared to the research on LLMs.
Figure 2. Statistical overview of research trends. (a) The number of papers on diffusion models at top computer vision conferences (CVPR, ECCV, ICCV) and top machine learning conferences (NeurIPS, ICML, ICLR) since 2017, indicating a growing interest. Note that ECCV and ICCV are held biennially. (b) Comparison of the ratio of papers on LLMs vs. diffusion models (left pie) and the research focus on alignment within these areas (right pie), highlighting the nascent stage of diffusion model alignment.
Inspired by the success of aligning LLMs (Wang et al., 2023d), there is growing interest in better aligning diffusion models with human intentions to enhance their capabilities. Fig. 2(b) visualizes the paper counts on LLMs and diffusion models, as well as their alignment studies  [^1] Data obtained from Google Scholar as of January 15, 2026. . The left pie chart shows LLMs account for 89.4% of the studies, while diffusion models account for 10.6%. The right chart highlights that, at this point, 93.5% of alignment studies focus on LLMs, while only 6.5% address diffusion models. This disparity, clearly illustrated in Fig. 2(b), not only underscores the relatively nascent stage of alignment research for diffusion models compared to LLMs but also signals a significant opportunity and pressing need for focused investigation in this domain. This survey aims to facilitate such research by consolidating current knowledge and outlining future directions. The fundamental differences between LLMs and diffusion models, where LLMs predict sequences of tokens and diffusion models progressively reverse a noise-adding diffusion process, and where the continuous, high-dimensional nature of image generation versus discrete token prediction in LLMs presents unique challenges for defining preferences and applying feedback, along with their uniquely advantageous application domains, such as LLMs for language generation and diffusion models for image generation, make the study of diffusion model alignment an independent and valuable area of interest.
@@ 3.1. Preference Data and Modeling
Preference data are typically composed of prompts, responses, and feedback. Prompts, whether human-provided or AI-generated, are diverse inputs used to elicit model responses. Responses can be generated on-policy (from current model) or off-policy (from external or older models), presenting a trade-off between relevance and collection efficiency (Sutton and Barto, 2020).
@@ 4. Human Alignment Techniques of Diffusion Models
RLHF and DPO are two very classic training-based techniques for aligning AI models with human preferences. However, when applied to diffusion models, these methods encounter significant challenges due to the step-by-step training and sampling nature of diffusion models. Specifically, aligning diffusion models with preference optimization requires sampling across all possible diffusion trajectories leading to $x_{0}$, which is intractable in practice. While the LLM response is treated as a single output, diffusion models’ multiple latent image representations of each step need to be calculated and stored, leading to unreasonable high memory consumption and low computation efficiency. This makes these methods impractical for large-scale diffusion models. To address the high computational overhead associated with adapting alignment techniques to diffusion models, researchers often formulate the denoising process as a multi-step Markov decision process (MDP). The proposed diffusion alignment methods need to directly optimize the expected reward of an image output or update the policy based on human preferences to approximately perform policy gradient guided by a reward model. This formulation enables parameter updates at each step of the denoising process based on human preferences, thereby circumventing the significant computational costs. Table 2 highlights the core differences among RLHF, DPO and test-time alignment techniques for diffusion models.
Table 2. Comparison of Human Alignment Paradigms for Diffusion Models.
Paradigm Compute Cost Feedback / Reward Scalability Key Limitations Best Use Cases
RLHF High: multi-step rollouts, trajectory storage, RL optimization Explicit: scalar reward via learned or heuristic reward models Low: limited by annotation cost and unstable RL training High variance, training instability, reward hacking, memory intensive When rich, well-defined rewards are easier to learn than the policy itself
DPO Moderate: avoids RL loops and explicit reward model training Implicit: relative preference via log-likelihood ratios Moderate: simpler pipeline but depends on high-quality preference pairs Sensitive to distribution shift; limited robustness outside preference data When preferences are available but reward modeling is unreliable or costly
Test-Time Alignment Low–Moderate: added inference-time optimization, no retraining Hybrid: heuristics or external (differentiable / black-box) rewards High: model-agnostic and training-free Increased inference latency; local or proxy-driven failures possible Lightweight, on-the-fly, or personalized alignment without model updates
@@ 4.3.3. Summary and Outlook
Test-time alignment techniques provide efficient and flexible ways to improve generation without retraining, evolving from implicit guidance to explicit reward-guided strategies. Implicit methods (e.g., prompt optimization or attention control) guide generation via heuristic or model-intrinsic modifications with minimal inference overhead, but offer limited control precision. In contrast, explicit methods incorporate external preference or reward models to directly optimize the sampling trajectory at inference time, enabling stronger alignment at the cost of higher computation and potential reward hacking.
Overall, the key distinction lies in whether an explicit preference objective is introduced at inference time: implicit methods modify inputs or internal states without an explicit objective, whereas explicit methods optimize generation with respect to an external preference signal. Control-input methods (e.g., spatial masks or sketches) are beyond the scope of this survey, as they specify what to generate rather than how preference alignment is enforced during sampling; see Cao et al. (2025) for a comprehensive survey of controllable generation.
@@ 4.5. Open Challenges and Research Frontiers in Diffusion Alignment
In this subsection, we discuss open challenges and research frontiers in diffusion alignment, focusing on the nuanced comparison between dominant paradigms, insights from LLM alignment, and challenges unique to the diffusion models.
@@ 5.1.3. AI Feedback Datasets
HPD v1 $\rightarrow$ HPS v1  (Wu et al., 2023b) Stable Foundation Discord channel 25,205 Stable Diffusion 98,807 2659 experienced users 25,205
HPD v2 $\rightarrow$ HPS v2  (Wu et al., 2023a) COCO Captions + DiffusionDB 107,915 9 models + COCO images 433,760 57 employed annotators 798,090
Pick-a-Pic v1 $\rightarrow$ PickScore  (Kirstain et al., 2023) Real users 37,523 Stable Diffusion variants 1,169,494 6,394 web app users 584,747
ImageRewardDB $\rightarrow$ ImageReward  (Xu et al., 2023) DiffusionDB 8,878 6 models 273,784 Annotation company 136,892
MHP $\rightarrow$ MPS  (Zhang et al., 2024b) PromptHero + DiffusionDB + KOLORS + GPT4 66,389 9 models 607,541 210 crowd-sourced annotators 918,315
RichHF-18K $\rightarrow$ RAHF  (Liang et al., 2024a) Pick-a-Pic v1 17,760 Pick-a-Pic v1 35,520 27 trained annotators 17,760
Picsart Image-Social $\rightarrow$ Social Reward  (Isajanyan et al., 2024) Social platform user prompts 104 K Several in-house models 1.7 M 1.5 M users 3M
VisionPrefer $\rightarrow$ VP-Score  (Wu et al., 2024c) DiffusionDB 179 K Stable Diffusion variants 0.76 M GPT-4 Vision 1.2 M
@@ 5.2.1. Evaluation for Reward Models
CLIP score (Radford et al., 2021) 54.3 71.2 63.7 60.8* 51.9
Aesthetic score (Schuhmann et al., 2022) 57.4 72.6 62.9 56.8* 55.3
HPS v1 (Wu et al., 2023b) 61.2 73.1 65.5 66.7* -
HPS v2 (Wu et al., 2023a) 65.7 83.3 65.5 67.4* 59.4
PickScore (Kirstain et al., 2023) 62.9 79.8 69.5 70.5* 62.6
ImageReward (Xu et al., 2023) 65.1 70.6 67.5 61.1* 60.5
MPS (Zhang et al., 2024b) 67.5 83.5 74.2 - -
VP-Score (Wu et al., 2024c) 66.3* 79.4* - 67.1* -
Social Reward (Isajanyan et al., 2024) - - - - 69.7
@@ 7. Summary
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No. 62506317.